### v2

# Neural Architecture Search via Two Constant Shared Weights Initialisations

Abstract
In recent years, zero-cost metrics are gaining ground in neural architecture search (NAS). These metrics allow finding the optimal neural network for a given task faster and with a lesser computational load than conventional NAS methods. Equally important is that they also shed some light on the internal workings of neural architectures. This paper presents a zero-cost metric that highly correlates with the train set accuracy across the NAS-Bench-101, NAS-Bench-201 and NAS-Bench-NLP benchmark datasets. We evaluate a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations. For this, we only use an unlabelled mini-batch of data. We observe that the dispersion of the outputs between two initialisations positively correlates with trained accuracy. The correlation further improves when we normalise dispersion by average output magnitude. The resulting metric, epsilon, does not require gradients computation and unbinds the NAS procedure from training hyperparameters, loss metrics and human-labelled data. Our method is easy to integrate within existing NAS algorithms and takes a fraction of a second to evaluate a single network. The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
Keywords: machine learning, neural architecture search, zero-cost NAS
1 Introduction
The field of neural architecture search (NAS) emerged about a decade ago as an effort to automatise the process of neural geometry optimisation. At the early stages of NAS development, every candidate architecture used to be evaluated through the training process (reinforcement learning (Williams (1992)), evolutionary algorithms (Real et al. (2019)), Bayesian optimisation (Falkner et al. (2018); White et al. (2021))).
One-shot algorithms adopting weight sharing dispense of multiple architectures training, reducing the search time drastically (efficient reinforcement learning (Pham et al. (2018)), random search with parameters sharing (Li and Talwalkar (2020)), differentiable methods). Nevertheless, they require the training of a massive hypernet, which necessitates elaborate hyperparameter tuning. While these methods prove efficient, they do not systematically achieve satisfactory results (Dong and Yang (2019)). One of the best of them, DARTS- (Chu et al. (2020)), shows significant uncertainty compared to evolutionary or reinforcement algorithms.
There are available methods that estimate network performance without training the dataset of interest but relying on an auxiliary predictive machine learning (ML) model built on a dataset of trained architectures (Istrate et al. (2019); Deng et al. (2017)). These methods aim to accelerate the NAS process for image recognition but still rely on training and cannot apply to other ML problems.
[…]
1.1 Zero-cost NAS
To alleviate the process of architecture search, many researchers focus on developing methods that allow finding optimal architectures without model training – so-called zero-cost NAS methods. These methods evaluate networks via a trainless proxy. They typically require the equivalent of one or a few training epochs, which is two to three orders of magnitude faster than previous NAS methods.
Weight agnostic neural networks. One of the pioneering works in zero-shot NAS is presented by Gaier and Ha (2019). The authors demonstrate a constructor building up neural architectures based on the mean accuracy over several initialisations with constant shared weights and the number of parameters contained within the model. The resulting model achieves over $90\%$ accuracy on MNIST data (LeCun et al. (2010)) when the weights are fixed to the best-performing constants. While these results are very intriguing, the authors admit that such architectures do not perform particularly well upon training. Moreover, back in $2019$, the benchmark databases of trained architectures, which are now routinely used to compare NAS metrics with each other, were yet to be released, which disables the comparison of this zero-shot method against the most recent ones.
ReLU activation pattern. In $2020$, Mellor et al. (2020) present the naswot metric, which exploits the rectified linear unit (ReLU, Agarap (2018)) activation function’s property to yield distinct activation patterns for different architectures. Concretely, every image yields a binary activation vector upon passing through a network, forming a binary matrix for a mini-batch. The logarithm of the determinant of this matrix serves as a scoring metric. Authors show that larger naswot values are associated with better training performances, which leads to the conclusion that high-performing networks should be able to distinguish the inputs before training. Unfortunately, the method can only be implemented on networks with ReLU activation functions, which limits its applicability to convolutional architectures.
Another work employing the aforesaid ReLU property is the one from Chen et al. (2021). They combine the number of linear regions in the input space with the spectrum of the neural tangent kernel (NTK) to build the tenas metric. Instead of evaluating each network in the search space individually, they create a super-network built with all the available edges and operators and then prune it.
Coefficient of variance. Our previous work on fully trainless NAS (Gracheva (2021)) evaluates the stability of untrained scores over random weights initialisations. The networks are initialised with multiple random seeds, and architectures are selected based on the coefficient of variance of the accuracy at initialisation, cv. While cv performance is associated with a high error rate, the author concludes that a good architecture should be stable against random weight fluctuations. While this method can, in theory, apply to any neural architecture type, it requires multiple initialisations and is relatively heavy compared to naswot and later methods. Furthermore, accuracy-based scoring metrics can only apply to classification problems, and it is unclear how to extend cv implementation to the regression tasks.
Gradient sign. The grad_sign metric is built to approximate the sample-wise optimisation landscape (Zhang and Jia (2021)). The authors argue that the closer local minima for various samples sit to each other, the higher the probability that the corresponding gradients will have the same sign. The number of samples yielding the same gradient sign approximates this probability. It allows to evaluate the smoothness of the optimisation landscape and architecture trainability. The method requires labels and gradient computation.
Pruning-at-initialisation proxies. Several powerful zero-cost proxies have emerged as an adaptation of pruning-at-initialisation methods to NAS in the work by Abdelfattah et al. (2021): grad_norm (Wang et al. (2020)), snip (Lee et al. (2018)), synflow (Tanaka et al. (2020)). These metrics are originally developed to evaluate the network’s parameters’ salience and prune away potentially meaningless synapses. They require a single forward-backwards pass to compute the loss. Then, the importance of parameters is computed as a multiplication of the weight value and gradient value. Then, they integrate the salience over all the parameters in the network to evaluate its potential upon training. What is particularly interesting about the synflow metric is that it evaluates the architectures without looking at the data by computing the loss as the product of all the weights’ values (randomly initialised). synflow metric shows the most consistent performance among various search spaces and sets the state-of-the-art for the zero-cost NAS.
Both naswot and synflow do not depend on labels, which arguably reduces the effect of human error during data labelling. Moreover, naswot does not require gradient computation, which renders this method less memory-intensive.
The above results imply that neural networks might have some intrinsic property which defines their prediction potential before training. Such property should not depend on the values of trainable parameters (weights) but only on the network’s topology. In the present work, we combine the takeaways from the existing trainless NAS implementations to present a new metric which significantly outperforms existing zero-cost NAS methods.
1.2 NAS benchmarks
To guarantee our metric’s reproducibility and compare its performance against other NAS algorithms, we evaluate it on the three widely used NAS benchmark datasets.
NAS-Bench-101. The first and one of the largest NAS benchmark sets of trained architectures. It consists of $423{,}624$ cell-based convolutional neural networks (Ying et al. (2019)). The architectures consist of three stacks of cells, each followed by max-pooling layers. Cells may have up to $7$ vertices and $9$ edges, with $3$ possible operations. This benchmark is trained multiple times on a single dataset, CIFAR-10 (Krizhevsky et al. (2009)), with a fixed set of hyperparameters for $108$ epochs.
NAS-Bench-201. It is a set of architectures with a fixed skeleton consisting of a convolution layer and three stacks of cells connected by a residual block (Dong and Yang (2020)). Each cell is a densely connected directed acyclic graph with $4$ nodes, $5$ possible operations and no limits on the number of edges, providing a total of $15{,}625$ possible architectures. Architectures are trained on three major datasets: CIFAR-10, CIFAR-100 (Krizhevsky et al. (2009)) and a downsampled version of ImageNet (Chrabaszcz et al. (2017)). Training hyperparameters are fixed, and the training spans $200$ epochs.
NAS-Bench-NLP. As the name suggests, this benchmark consists of architectures suitable for neural language processing (Klyuchnikov et al. (2022)). Concretely, it consists of randomly-generated recurrent neural networks. Recurrent cells comprise $24$ nodes, $3$ hidden states and $3$ input vectors at most, with $7$ allowed operations. Here, we only consider models trained and evaluated on Penn Treebank (PTB, Marcinkiewicz (1994)) dataset: $14{,}322$ random networks with a single seed and $4{,}114$ with tree seeds. The training spans $50$ epochs and is conducted with fixed hyperparameters.
2 Epsilon metric
Two existing NAS methods inspire the metric that we share in the present work: cv (Gracheva (2021)) and weight agnostic neural networks (Gaier and Ha (2019)). Both metrics aim to exclude individual weight values from consideration when evaluating networks: the former cancels out the individual weights via multiple random initialisations, while the latter sets them to the same value across the network. It is very intriguing to see that an network can be characterised purely by its topology.
As mentioned above, the cv metric has two principal disadvantages. While it shows a fairly consistent trend with trained accuracy, it suffers from high uncertainty. It can be, to some degree, explained by the fact that random weight initialisations bring in some noise. Our idea is that replacing random initialisations with single shared weight initialisations, similarly to Gaier and Ha (2019), should improve the method’s performance.
The second weak point is that cv is developed for classification problems and relies on accuracy. Therefore, it needs to be clarified how to apply this metric to regression problems. The coefficient of variation is a ratio of standard deviation to mean, and Gracheva (2021) shows that cv correlates negatively with train accuracy. It implies that the mean untrained accuracy should be maximised. On the other hand, for regression tasks, performance is typically computed as some error, which is sought to be minimised. It is not apparent whether the division by mean untrained error would result in the same trend for cv metric.
To address this issue, we decided to consider raw outputs. This modification renders the method applicable to any neural architecture. However, it comes with a difference: accuracy returns a single value per batch of data, while raw outputs are $[N_{\textrm{BS}}\times L]$ matrices, where $N_{\rm{BS}}$ is the batch size and $L$ is the length of a single output. [^1] This length depends on the task and architecture: for regression tasks $L=1$, for classification, $L$ is equal to the number of classes, and for recurrent networks, it depends on the desired length of generated string. In our work, we flatten these matrices to obtain a single vector $\mathbf{v}$ of length $L_{v}=N_{\textrm{BS}}\times L$ per initialisation. We then stuck both initialisations into a single output matrix $\mathbf{V}$.
Before proceeding to statistics computation over initialisations, we also must normalise the output vectors: in the case of constant shared weights, outputs scale with weight values. In order to compare initialisations on par with each other, we use min-max normalisation:
$\mathbf{V}_{i}^{\prime}=\frac{\mathbf{V}_{i}-\min(\mathbf{V}_{i})}{\max(\mathbf{V}_{i})-\min(\mathbf{V}_{i})},$ (1)
where $i$ is the index for initialisations, $i=\{0,1\}$.
We noticed that two distinct weights are sufficient to grasp the difference between initialisations. Accordingly, instead of standard deviation, we use mean absolute error between the normalised outputs of two initialisations:
$\delta=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\lvert\mathbf{V}_{1,j}^{\prime}-\mathbf{V}_{2,j}^{\prime}\rvert.$ (2)
The mean is computed over the outputs of both initialisations as follows:
$\mu=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\frac{\mathbf{V}_{1,j}^{\prime}+\mathbf{V}_{2,j}^{\prime}}{2}=\frac{1}{2L_{v}}\sum\limits_{i=0}^{2}\sum\limits_{j=0}^{L_{v}}\mathbf{V}_{i,j}^{\prime}$ (3)
Finally, the metric is computed as the ratio of $\delta$ and $\mu$:
$\varepsilon=\frac{\delta}{\mu}.$ (4)
We refer to our metric as epsilon, as a tribute to the $\varepsilon$ symbol used in mathematics to denote error bounds. Algorithm 1 details the epsilon metric computation.
Algorithm 1 Algorithm for epsilon metric computation
  Select a batch of data from train set
  for arch in search space do
     Initialise empty output matrix
     for weight in [val1, val2do
        Initialise arch with constant shared weight
        Forward pass the batch through arch
        Get and flatten outputs
        Minmax normalise outputs (Eq. 1)
        Append outputs to the output matrix
     end for
     Compute difference between the rows of the output matrix (Eq. 2)
     Compute mean over the output matrix (Eq. 3)
     Compute epsilon metric (Equation 4)
  end for
2.1 Empirical evaluation
Here we evaluate the performance of epsilon and compare it to the results for zero-cost NAS metrics reported in Abdelfattah et al. (2021). We use the following evaluation scores (computed with NaN omitted):
  • Spearman $\rho$ (global): Spearman rank correlation $\rho$ evaluated on the entire dataset.
  • Spearman $\rho$ (top-$10\%$): Spearman rank correlation $\rho$ for the top-$10\%$ performing architectures.
  • Kendall $\tau$ (global): Kendall rank correlation coefficient $\tau$ evaluated on the entire dataset.
  • Kendall $\tau$ (top-$10\%$): Kendall rank correlation coefficient $\tau$ for the top-$10\%$ performing architectures.
[…]
  • Top-64/top-$10\%$: number of top-64 models ranked by zero-cost scoring metric within top-$5\%$ performing models.
2.1.1 NAS-Bench-201
The results for overall epsilon performance on NAS-Bench-201 are given in Table 1 along with other zero-cost NAS metrics. The Kendall $\tau$ score is not reported in Abdelfattah et al. (2021), but it is considered more robust than Spearman $\rho$ and is increasingly used for NAS metric evaluation. We use the data provided by Abdelfattah et al. (2021) to evaluate their Kendal $\tau$. Note that our results differ from the original paper for some evaluation scores. In such cases, we indicate the original values between brackets. In particular, there is a discrepancy in computing the values in the last column, Top-64/top-$5\%$, while the rest of the results are consistent. Figure 6 in Appendix suggests that our calculations are correct.
Table 1: Zero-cost metrics performance for the NAS-Bench-201 search space with its three datasets: CIFAR-10, CIFAR-100 and ImageNet16-120. We give the original values from Abdelfattah et al. (2021) for reference between brackets. We highlight the best-performing metrics in bold.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
CIFAR-10
grad_sign 0.77
synflow 0.74 0.18 0.54 0.12 45.75 (46) 29 (44)
grad_norm 0.59 (0.58) -0.36 (-0.38) 0.43 -0.21 30.26 (30) 1 (0)
grasp 0.51 (0.48) -0.35 (-0.37) 0.36 -0.21 30.77 (30) 3 (0)
snip 0.60 (0.58) -0.36 (-0.38) 0.44 -0.21 30.65 (31) 1 (0)
fisher 0.36 -0.38 0.26 -0.24 4.99 ( 5) 0 (0)
epsilon 0.87 0.55 0.70 0.40 67.39 59
CIFAR-100
grad_sign 0.79
synflow 0.76 0.42 0.57 0.29 49.71 (50) 45 (54)
grad_norm 0.64 -0.09 0.47 -0.05 35.00 (35) 0 (4)
grasp 0.55 (0.54) -0.10 (-0.11) 0.39 -0.06 35.32 (34) 3 (4)
snip 0.64 (0.63) -0.08 (-0.09) 0.47 -0.05 35.25 (36) 0 (4)
fisher 0.39 -0.15 (-0.16) 0.28 -0.10 4.22 ( 4) 0 (0)
epsilon 0.90 0.59 0.72 0.43 81.24 62
ImageNet16-120
grad_sign 0.78
synflow 0.75 0.55 0.56 0.39 43.57 (44) 26 (56)
grad_norm 0.58 0.12 (0.13) 0.43 0.09 31.29 (31) 0 (13)
grasp 0.55 (0.56) 0.10 0.39 0.07 31.61 (32) 2 (14)
snip 0.58 0.13 0.43 0.09 31.16 (31) 0 (13)
fisher 0.33 0.02 0.25 0.01 4.61 ( 5) 0 (0)
epsilon 0.85 0.53 0.67 0.37 71.51 59
For NAS-Bench-201, we also report average performance when selecting one architecture from a pool of $N$ random architectures. The statistics are reported over $500$ runs. Table 2 compares epsilon to other trainless metrics. Note that te-nas starts with a super-network composed of all the edges and operators available within the space. In this case, $N$ is not applicable, and the performance can not be improved. In principle, other methods’ performance should improve with higher $N$ values.
Table 2: Comparison of the trainless metrics performances against existing NAS algorithms on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. On the top, we list the best-performing methods that require training: REA (Real et al. (2019)), random search, REINFORCE (Williams (1992)), BOHB (Falkner et al. (2018)), DARTS- (Chu et al. (2020)). We report the average best-achieved test accuracy over $500$ runs, with $1{,}000$ architectures ($100$ for grad_sign) sampled from the search space at random. For tenas, the results are reported for $4$ random seeds. Random and optimal performances are given as a baseline.
Method Cost CIFAR-10 CIFAR-100 ImageNet16-120
(h) validation test validation test validation test
State-of-the-art
REA $3.3$ $91.19\pm 0.31$ $93.92\pm 0.3$ $71.81\pm 1.12$ $71.84\pm 0.99$ $45.15\pm 0.89$ $45.54\pm 1.03$
RS $3.3$ $90.93\pm 0.36$ $93.92\pm 0.31$ $70.93\pm 1.09$ $71.04\pm 1.07$ $44.45\pm 1.1$ $44.57\pm 1.25$
REINFORCE $3.3$ $91.09\pm 0.37$ $93.92\pm 0.32$ $71.61\pm 1.12$ $71.71\pm 1.09$ $45.05\pm 1.02$ $45.24\pm 1.18$
BOHB $3.3$ $90.82\pm 0.53$ $93.92\pm 0.33$ $70.74\pm 1.29$ $70.85\pm 1.28$ $44.26\pm 1.36$ $44.42\pm 1.49$
DARTS- $3.2$ $91.03\pm 0.44$ $93.80\pm 0.40$ $71.36\pm 1.51$ $71.53\pm 1.51$ $44.87\pm 1.46$ $45.12\pm 0.82$
Baselines (N=1000)
Optimal n/a $91.34\pm 0.18$ $94.20\pm 0.13$ $72.53\pm 0.53$ $72.84\pm 0.41$ $45.93\pm 0.51$ $46.59\pm 0.34$
Random n/a $84.11\pm 11.71$ $87.40\pm 11.94$ $61.57\pm 11.35$ $61.67\pm 11.35$ $33.97\pm 8.68$ $33.67\pm 8.98$
Trainless (N=1000)
naswot $3.3$ $89.69\pm 0.73$ $92.96\pm 0.81$ $69.86\pm 1.21$ $69.98\pm 1.22$ $43.95\pm 2.05$ $44.44\pm 2.10$
synflow n/a $89.91\pm 0.83$ $90.12\pm 0.78$ $70.35\pm 2.25$ $70.37\pm 2.08$ $41.73\pm 3.91$ $42.11\pm 4.02$
grad_norm n/a $88.13\pm 2.35$ $88.42\pm 2.28$ $66.35\pm 5.45$ $66.48\pm 5.32$ $33.88\pm 11.46$ $33.90\pm 11.74$
grasp n/a $87.85\pm 2.12$ $88.17\pm 2.04$ $65.36\pm 5.57$ $65.45\pm 5.48$ $32.23\pm 10.95$ $32.20\pm 11.23$
snip n/a $87.47\pm 2.19$ $87.81\pm 2.12$ $64.61\pm 5.52$ $64.74\pm 5.43$ $30.65\pm 11.32$ $30.55\pm 11.55$
fisher n/a $87.01\pm 2.31$ $87.36\pm 2.23$ $63.54\pm 5.69$ $63.67\pm 5.62$ $26.70\pm 10.83$ $29.56\pm 10.83$
epsilon $0.03$ $\boldsymbol{91.03}\pm\boldsymbol{0.42}$ $\boldsymbol{93.86}\pm\boldsymbol{0.43}$ $\boldsymbol{71.76}\pm\boldsymbol{0.90}$ $\boldsymbol{71.79}\pm\boldsymbol{0.86}$ $\boldsymbol{45.11}\pm\boldsymbol{0.99}$ $\boldsymbol{45.42}\pm\boldsymbol{1.21}$
Trainless (N=100)
cv $0.05$ $84.89\pm 6.39$ $91.90\pm 2.27$ $63.99\pm 5.61$ $64.08\pm 5.63$ $38.68\pm 6.34$ $38.76\pm 6.62$
grad_sign n/a $89.84\pm 0.61$ $93.31\pm 0.47$ $70.22\pm 1.32$ $70.33\pm 1.28$ $42.07\pm 2.78$ $42.42\pm 2.81$
epsilon $0.003$ $\boldsymbol{90.44}\pm\boldsymbol{0.97}$ $\boldsymbol{93.39}\pm\boldsymbol{0.82}$ $\boldsymbol{70.85}\pm\boldsymbol{1.30}$ $\boldsymbol{71.00}\pm\boldsymbol{1.26}$ $\boldsymbol{44.03}\pm\boldsymbol{2.02}$ $\boldsymbol{44.20}\pm\boldsymbol{2.04}$
Trainless (all space)
tenas $0.5$ $93.9\pm 0.47$ $71.24\pm 0.56$ $42.38\pm 0.46$
Comparing the results for epsilon with other zero-cost NAS metrics, we can see that it outperforms them by a good margin. Figure 1 further confirms the applicability of the method to the NAS-Bench-201 field (similar figures for other methods can be found in Appendix, Figure 6). However, NAS-Bench-201 is a relatively compact search space; furthermore, it has been used for epsilon development.
Refer to captionRefer to captionRefer to caption
Figure 1: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure represents the search space of $15{,}625$ networks (excluding architectures with NaN scores).
2.1.2 NAS-Bench-101
We use NAS-Bench-101 space to confirm that the success of epsilon metric in the previous section is not due to overfitting the NAS-Bench-201 search space and to see how it applies to a vaster search space. Table 3 together with Figure 2 confirm that it performs reasonably well on NAS-Bench-101, too.
Table 3: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. Values from Abdelfattah et al. (2021) are given for reference between brackets. We highlight the best-performing metrics in bold.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
grad_sign 0.45
synflow 0.37 0.14 0.25 0.10 22.67 (23) 4 (12)
grad_norm -0.20 -0.05 (0.05) -0.14 -0.03 1.98 (2) 0 (0)
grasp 0.45 -0.01 0.31 -0.01 25.60 (26) 0 (6)
snip -0.16 0.01 (-0.01) -0.11 0.00 3.43 (3) 0 (0)
fisher -0.26 -0.07 (0.07) -0.18 -0.05 2.65 (3) 0 (0)
epsilon 0.62 0.12 0.44 0.08 40.33 10
Refer to captionRefer to caption
Figure 2: Zero-cost NAS epsilon metric performance illustration for NAS-Bench-101 search space, CIFAR-10 dataset and NAS-Bench-NLP search space, PTB dataset. The horizontal axis shows test accuracy upon training. Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. The figure shows $423{,}624$ and $14{,}322$ networks for NAS-Bench-101 and NAS-Bench-NLP, respectively (excluding architectures with NaN scores).
2.1.3 NAS-Bench-NLP
Both NAS-Bench-201 and NAS-Bench-101 are created to facilitate NAS in image recognition. They operate convolutional networks of very similar constitutions. To truly probe the generalisability of the epsilon metric, we test it on NAS-Bench-NLP. Both input data format and architecture type differ from the first two search spaces.
Unfortunately, Abdelfattah et al. (2021) provides no data for NAS-Bench-NLP, disabling us from using their results for calculations. Therefore, in Table 4, we give only values provided in the paper together with our epsilon metric (data for ficher is absent). We want to note that, unlike accuracy, perplexity used for language-related ML problems should be minimised. Therefore, the signs of correlations with scoring metrics should be reversed, which is not the case for numbers given in Abdelfattah et al. (2021).
Table 4: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We highlight the best-performing metrics in bold.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
synflow 0.34 0.10 22
grad_norm -0.21 0.03 10
grasp 0.16 0.55 4
snip -0.19 -0.02 10
epsilon -0.34 -0.12 -0.23 -0.08 24.87 11
2.2 Integration with other NAS methods
While it is possible to utilise zero-cost metrics independently, they are often implemented within other NAS. Here we provide examples of random search and ageing evolution algorithms when used in tandem with epsilon.
Similarly to Abdelfattah et al. (2021), we compare random search performance with and without warm-up. First, we create a warm-up pool of $3{,}000$ architectures. Then, during the first $64$ steps of random search, the algorithm picks networks from the pool based on the highest trainless score and accordingly updates the best test performance.
For ageing evolution, the same principle applies. En plus, we report the results of implementation where every next parent is decided based on the highest epsilon score (”move”). In other words, the trainless scoring metric replaces validation accuracy in move mode. Finally, the child is created by parent mutation (within an edit distance of $1$) and added to the pool. Finally, the oldest network is removed from the pool.
For both described algorithms, we run the procedure until the number of trained architectures reaches $300$ and perform $100$ random rounds. Figure 3 shows that epsilon metric leads to considerable improvements in terms of time and precision. The best performance is achieved in combination with a warm-up. Figure 5 assembles warm-up performances for several trainless metrics.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: epsilon integration within ageing evolution (top) and random search (bottom) NAS algorithms for three datasets from NAS-Bench-201 search space.
3 Ablation studies
While our metric is relatively straightforward to implement, it does require several hyperparameters to be set. In this section, we provide the ablations studies’ results and analyse them.
3.1 Weights
To compute the epsilon metric, we initialise the networks with two distinct constant shared weights. The question is: how to fix their values, and how does this choice affect the whole method?
To answer this question, we ran a series of tests with multiple pairs of weights. In pairs, it makes sense to set the first weight always greater than the second one. This consideration stems from the fact that epsilon is based on the mean absolute difference between two initialisations. Therefore, it will output the same score for [$w_{1}$, $w_{2}$] and [$w_{2}$, $w_{1}$] combinations and will be exactly zero in the case where the two weights are equal. Tables 5, 6, 7 summarises the results of our tests. For the NAS-Bench-201 search space, the outcomes are very similar between the datasets, and the minor differences do not affect the conclusions drawn from this data.
Table 5: Zero-cost metrics performance evaluated on NAS-Bench-101 search space, CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
0.0001, 0.001 2120 0.47 0.41 0.33 0.28 38.21 23
0.0001, 0.01 2120 0.61 0.03 0.43 0.02 41.98 24
0.0001, 0.1 2120 0.61 0.03 0.43 0.02 41.98 24
0.0001, 1 2120 0.61 0.03 0.43 0.02 41.98 24
0.001, 0.01 4968 0.29 -0.05 0.20 -0.03 17.30 10
0.001, 0.1 4968 0.29 -0.05 0.19 -0.03 17.30 8
0.001, 1 4968 0.29 -0.05 0.19 -0.03 17.30 8
0.01, 0.1 5000 0.21 -0.12 0.14 -0.08 4.60 0
0.01, 1 5000 0.21 -0.12 0.14 -0.08 4.60 0
0.1, 1 5000 -0.47 0.04 -0.32 0.03 0.20 0
Table 6: Zero-cost metrics performance evaluated on NAS-Bench-201 search space, CIFAR-10 dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
1e-07, 1e-06 2957 0.33 -0.20 0.22 -0.13 4.76 1
1e-07, 1e-05 2957 0.39 -0.06 0.29 -0.04 9.15 2
1e-07, 0.0001 2957 0.47 -0.01 0.32 0.00 9.12 2
1e-07, 0.001 2957 0.62 0.06 0.43 0.05 18.92 3
1e-07, 0.01 2957 0.83 0.69 0.64 0.49 68.40 56
1e-07, 0.1 2957 0.84 0.64 0.65 0.46 68.03 56
1e-07, 1 2957 0.87 0.59 0.70 0.43 65.88 45
1e-06, 1e-05 2957 0.41 -0.06 0.30 -0.04 9.15 2
1e-06, 0.0001 2957 0.47 -0.01 0.32 0.00 9.12 2
1e-06, 0.001 2957 0.62 0.06 0.43 0.05 18.92 3
1e-06, 0.01 2957 0.83 0.69 0.64 0.49 68.40 56
1e-06, 0.1 2957 0.84 0.64 0.65 0.46 68.03 56
1e-06, 1 2957 0.87 0.59 0.70 0.43 65.88 45
1e-05, 0.0001 2957 0.47 -0.01 0.32 0.00 9.12 2
1e-05, 0.001 2957 0.62 0.06 0.43 0.05 18.92 3
1e-05, 0.01 2957 0.83 0.69 0.64 0.49 68.40 56
1e-05, 0.1 2957 0.84 0.64 0.65 0.46 68.03 56
1e-05, 1 2957 0.87 0.59 0.70 0.43 65.88 45
0.0001, 0.001 2957 0.62 0.06 0.43 0.05 19.26 3
0.0001, 0.01 2957 0.82 0.69 0.63 0.49 68.42 56
0.0001, 0.1 2957 0.84 0.64 0.65 0.46 68.14 56
0.0001, 1 2957 0.87 0.58 0.69 0.42 65.88 46
0.001, 0.01 4802 0.25 0.41 0.17 0.28 30.35 52
0.001, 0.1 4802 0.31 0.27 0.20 0.19 26.82 25
0.001, 1 4802 0.31 0.22 0.21 0.15 26.26 6
0.01, 0.1 4907 0.15 0.42 0.10 0.29 37.47 57
0.01, 1 4907 0.12 0.39 0.06 0.27 35.85 57
0.1, 1 4907 0.01 0.26 -0.03 0.16 0.84 0
Table 7: Zero-cost metrics performance evaluated on NAS-Bench-NLP search space, PTB dataset. We start with $5{,}000$ randomly picked architectures. The number of architectures with non-NaN values are given in the second column.
Weights Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
1e-07, 1e-06 2393 -0.24 0.32 -0.16 0.22 0.00 0
1e-07, 1e-05 2391 -0.22 0.52 -0.15 0.39 1.67 0
1e-07, 0.0001 2388 -0.35 0.49 -0.24 0.37 0.42 1
1e-07, 0.001 2332 -0.42 0.50 -0.29 0.37 2.99 1
1e-07, 0.01 1215 -0.30 0.13 -0.21 0.09 4.20 1
1e-07, 0.1 585 -0.49 0.55 -0.36 0.44 1.69 1
1e-07, 1 403 -0.12 0.08 -0.08 -0.00 12.20 7
1e-06, 1e-05 3286 -0.11 0.49 -0.07 0.36 0.91 0
1e-06, 0.0001 3283 -0.24 0.47 -0.16 0.34 0.62 0
1e-06, 0.001 3226 -0.33 0.44 -0.23 0.33 4.64 1
1e-06, 0.01 1639 -0.26 0.13 -0.18 0.09 3.40 2
1e-06, 0.1 841 -0.41 0.13 -0.28 0.12 10.71 1
1e-06, 1 594 -0.09 0.13 -0.06 0.08 11.67 7
1e-05, 0.0001 3948 -0.32 0.30 -0.21 0.22 0.51 0
1e-05, 0.001 3892 -0.35 0.20 -0.24 0.16 3.85 2
1e-05, 0.01 1900 -0.32 0.27 -0.22 0.18 2.86 1
1e-05, 0.1 913 -0.38 0.31 -0.28 0.25 13.04 1
1e-05, 1 649 -0.18 0.36 -0.12 0.25 10.77 7
0.0001, 0.001 3987 -0.35 0.18 -0.24 0.14 3.76 0
0.0001, 0.01 1943 -0.37 0.38 -0.26 0.25 3.59 1
0.0001, 0.1 925 -0.43 0.26 -0.30 0.21 11.83 1
0.0001, 1 663 -0.26 0.35 -0.18 0.24 10.45 7
0.001, 0.01 1936 -0.32 0.40 -0.21 0.26 2.59 3
0.001, 0.1 918 -0.38 0.27 -0.30 0.22 8.70 2
0.001, 1 663 -0.23 -0.20 -0.19 -0.17 7.46 5
0.01, 0.1 910 -0.36 0.04 -0.25 0.02 8.79 2
0.01, 1 654 -0.25 0.06 -0.16 0.04 9.09 5
0.1, 1 652 -0.27 0.20 -0.19 0.13 10.61 7
The results show that there is indeed some variation between different initialisations. Two effects impact performance. Firstly, the more significant the difference between the two weights, the higher the correlation between epsilon and trained performance. It makes sense since if the weights are too close to each other, the difference between the outputs becomes minimal, and the distinction between architectures in terms of epsilon is no longer possible.
Secondly, we can see that too low and too high weights result in fewer architectures with non-NaN scores. Extreme weights result in all zeros or infinite output values and, consequently, in NaN value of the epsilon metric. Naturally, the deeper the network is, the more signal attenuation is, and the higher the probability of NaN score. We remind that prior to epsilon computation, the outputs are normalised to the same range $[0,1]$, which means the value of the metric per se does not become smaller for deeper networks. This extinction effect is due to the limited sensitivity of float type. Theoretically, with infinite float sensitivity, there would be no extinction and hence no effects on the choice of the weights.
Table 8: Optimal weights for epsilon evaluation for three search spaces and their datasets.
Search space Dataset Optimal weights
NAS-Bench-101 CIFAR-10 $[10^{-4},10]$
NAS-Bench-201 CIFAR-10 $[10^{-7},1]$
CIFAR-100 $[10^{-7},1]$
ImageNet16-120 $[10^{-7},1]$
NAS-Bench-NLP PTB $[10^{-5},10^{-3}]$
As a rule of thumb, for every new ML problem, we suggest running epsilon evaluation on a subset of architectures with several weights and selecting the minimum and maximum weights that do not cause excessive NaN outputs. While we acknowledge that this procedure is subjective, importantly, it does not require any training.
Weight combinations that we selected for each search space are given in Table 8.
3.2 Initialisation algorithm
The choice to initialise the weights with constant values provides a ground for discussion. On the one hand, constant shared weights guarantee the same initialisation regardless of the structure; the noise from the randomness of other initialisations is not obstructing the effect of the neural geometry. It would be possible to reduce the amount of noise by increasing the number of independent initialisations with distinct random seeds similar to Gracheva (2021), but this increases the algorithm run-time proportionally.
Table 9 gives the results of the ablation studies when we compute epsilon based on two initialisations.
Table 9: Initialisations ablation studies. The metric is computed over $5{,}000$ initial architectures. The number of remaining architectures is given in the second column.
Initialisation Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
Uniform 15284 0.14 -0.17 0.09 -0.11 3.07 7
Normal 15284 0.44 -0.29 0.30 -0.20 20.07 4
Kaiming uniform 15284 0.32 0.11 0.21 0.07 27.40 26
Kaiming normal 15284 0.49 -0.29 0.33 -0.20 17.86 7
Orthogonal 15284 0.41 -0.20 0.28 -0.13 15.11 1
3.3 Batch size
To test the method’s sensitivity against the batch size, we ran the scoring routine with $8$ different batch sizes $[8,16,32,64,128,512,1024]$, each with $10$ random batches. The tests are done on the NAS-Bench-201 search space with its three datasets. For each batch, we report medians with $25$ and $75$ percentiles over $10$ runs. We evaluate epsilon on $5{,}000$ architectures.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Batch size ablation study for NAS-Bench-201 search space, CIFAR-10 (left), CIFAR-100 (centre) and ImageNet16-120 (right) datasets.
Figure 4 show that, as expected, the larger the batch size, the better and more stable the performance of epsilon metric. There is almost no improvement for batches over $256$. We set the batch size to this value in the experiments throughout the paper.
3.4 Embedding initialisation
In order to verify the effect of the embedding initialisation, we have run tests with $6$ different initialisations:
  • uniform positive 0.1: random uniform from ranges [0, 0.1]
  • uniform positive 1: random uniform from ranges [0, 1]
  • uniform centred 0.1: random uniform from ranges [-0.1, 0.1]
  • uniform centred 1: random uniform from ranges [-1, 1]
  • random 0.1: random normal centred at 00 with a standard deviation of 0.1
  • random 1: random normal centred at 00 with a standard deviation of 1
Table 10 summarises our results of the embedding ablations. Ablation is done on the NAS-Bench-NLP search space, PTB dataset (the only search space implementing embedding).
Table 10: Embedding ablation studies. Metric computed over $5{,}000$ initial architectures. The number of remaining architectures is given in the second column.
Embedding Archs Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
Uniform positive 0.1 782 -0.38 0.17 -0.26 0.15 3.80 2
Uniform positive 1 776 -0.37 0.21 -0.26 0.18 2.56 2
Uniform centered 0.1 783 -0.40 0.18 -0.28 0.15 2.53 2
Uniform centered 1 782 -0.46 0.18 -0.33 0.16 1.27 1
Random 0.1 783 -0.42 0.18 -0.29 0.15 2.53 2
Random 1 782 -0.47 0.19 -0.33 0.16 1.27 1
Care should be taken when initialising the networks containing the embedding layer: if embedding is initialised with all constants, there is no difference between the embedded input. The performance of our metric, in this case, will be analogous to that with a batch size of one.
Our ablation studies show that as soon as the embedding is initialised with non-constant weights, it does not significantly influence the outcomes.
3.5 Synthetic data
In this section, we test the importance of the input data by feeding our metric several synthetic input data on CIFAR-10. We feed networks with $5$ types of data:
  • actual data: a batch of CIFAR-10 data
  • grey scale images: images within the batch are solid colour ranging from black to white
  • random normal: images are filled with random values following normal distribution with $[\mu,\sigma]=[0,1]$
  • random uniform: images are filled with random values following uniform distribution with $[\mu,\sigma]=[-1,1]$
  • random uniform (+): images are filled with random values following uniform distribution with $[min,\max]=[0,1]$
All the tests are performed with batch size of $256$, weights $[10^{-7},1]$ and $5{,}000$ architectures. Table 11 shows that even though the performance with synthetic data drops compared to actual data, it is still reasonably good.
Curiously, greyscale images, filled with constant values, show the closest results to the actual data. Note that epsilon metric with greyscale data outperforms synflow. It is an essential achievement because synflow does not use input data and is, therefore, data independent. Our results show that epsilon has the potential to be used with no data whatsoever.
Table 11: Synthetic data tests with CIFAR-10 dataset, NAS-Bench-201 search space. Tests are based on the evaluation of $5{,}000$ architectures.
Metric Spearman $\rho$ Kendall $\tau$ Top-10%/ Top-64/
global top-10% global top-10% top-10% top-5%
Real data 0.87 0.59 0.70 0.43 65.88 $45$
Grey scale 0.87 0.44 0.68 0.31 59.46 $34$
Random normal 0.54 0.15 0.38 0.14 14.97 $3$
Random uniform 0.56 0.17 0.40 0.16 14.19 $3$
Random uniform (+) 0.61 0.17 0.43 0.16 16.22 $4$
4 Visualisation of other zero-cost NAS metrics
4.1 When integrated with other NAS methods
In Section 2.2, we show how epsilon metric improves the performance of ageing evolution and random search when used for warming up. Figure 5 compares epsilon integration to other metrics from Abdelfattah et al. (2021). For both EA ans RS, we use the metrics for a warm-up and run the procedure until the number of trained architectures reaches $300$ with $100$ random rounds. The warm-up pool contains $3000$ randomly selected architectures.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Various zero-cost NAS metrics’ performance when used as a warm-up for ageing evolution (top) and random search (bottom) for three datasets from the NAS-Bench-201 search space. Shadows span between lower and upper quartiles based on $100$ rounds. The warm-up population is $3000$ architectures.
4.2 Correlation with accuracy
In Abdelfattah et al. (2021) the metrics are presented through statistical measures, but we feel that visualisation helps to improve understanding. Here we provide visualisations for two search spaces built on the data provided by the authors.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Zero-cost metrics performance illustration for NAS-Bench-201 search space evaluated on CIFAR-10, CIFAR-100 and ImageNet16-120 datasets, based on data from Abdelfattah et al. (2021). Each dot corresponds to an architecture; the darker the colour, the more parameters it contains. Figure represents search space of $15{,}625$ networks (excluding architectures with NaN scores).
5 Discussion
The epsilon metric performs surprisingly well on various benchmark datasets. Similarly to other NAS algorithms, it shows the most robust performance with the NAS-Bench-201 benchmark. This fact indicates that the set of architectures comprising NAS-Bench-201 forms a relatively easy-to-navigate landscape regardless of the metric. Still, epsilon shows the best performance among trainless metrics. Table 1 shows that synflow shows comparable performance in terms of correlations on the ImageNet-16-120 dataset. However, statistically epsilon selects, on average, a better-performing architecture, which can also be confirmed visually by comparing Figures 1 and 6.
Notably, we would like to highlight the difference with our previous work (cv metric, Gracheva (2021)). The two methods differ in how the weights are de-emphasised and what outputs we use for statistics computation. Two constant shared weight initialisations better bring out the effect of the bare geometry compared to a hundred random initialisations. Moreover, epsilon metric takes significantly less time to compute, being fifty-fold faster.
It is important to note that we used NAS-Bench-201 for the metric development; therefore, the results of this part of the study might be biased.
To confirm the generalisability of the metric, we applied it to the NAS-Bench-101 benchmark. Again, Table 3 shows strong performance. The peculiar vertical lines are due to clustering in terms of test accuracy. Since every architecture in NAS-Bench-101 is trained with $3$ random seeds, the four clusters correspond to situations when the training diverges for some random initialisations. More random seeds would reveal a smoother picture and better correlations. Again, we observe comparable performance between synflow and epsilon scoring metrics. We could only confirm some of the results from Abdelfattah et al. (2021) and suppose that some statistics might have been miscalculated.
By design, we expect our metric to be independent of the specific neural architecture structure. To confirm this claim, this is essential to show its applicability to architectures other than CNNs. Our work tests epsilon applicability to RNN models with NAS-Bench-NLP.
While statistics in Table 4 are not exceptional, Figure 2 shows a visible trend towards better architectures with increasing metric value. Significant noise levels might stem from the characteristics of the benchmark itself: factors like a relatively small sample of networks given a vast space, choice of training hyperparameters, dropout rates, and others may distort NAS performance. Nonetheless, the trend is strong enough to conclude that epsilon metric can apply to recurrent type architectures.
While epsilon metric shows solid empirical performance, the underlying reasons for this are unclear.
There are several hints towards its understanding. First, mathematically, epsilon represents the difference in the output distribution shapes between initialisations. The output shape is affected by layer widths, activation functions, batch normalisation, skip connections and other factors, which we generally call network geometry. By setting weights to a constant shared value, one can probe the effects of the geometry without being obstructed by the randomness of initialisation.
Second, according to the weight ablation studies (Section 3), the best performance is achieved when we set the weights to the lowest and highest values that do not lead to excessive outputs explosion or vanishing. Therefore, epsilon measures the amplitude of the outputs’ distribution shape change due to geometry.
Finally, we see that grey-scale solid images work reasonably well as inputs during the synthetic data studies. The distribution over the input samples is uniform, which makes it easier to track the changes in the distribution shape as the signal propagates through the network.
6 Conclusion
This work presents a novel zero-cost NAS scoring metric epsilon. It allows evaluating a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations. The value of epsilon reflects how neural network outputs’ distribution changes between low and high constant shared weight initialisations. It shows that the higher this difference, the better the network will perform upon training.
It takes a fraction of a second to evaluate a single architecture and does not require gradients computation (low demand on memory) or labels, and can run on a CPU. We provide experimental evidence that the introduced epsilon metric generalises to various architecture types, such as convolutional or recurrent neural networks (care should be taken with embedding layer initialisation).
This work evaluates epsilon on three staple NAS search spaces: NAS-Bench-201, NAS-Bench-101 and NAS-Bench-NLP. It sets the state-of-the-art among zero-cost NAS metrics and can be easily implemented within other methods, such as random search or evolutionary algorithms.
The only significant disadvantage of the method is that it requires the choice of constant weight values used for initialisation. Our tests show that it must be set up individually for each search space. We plan to automate the weight selection process in our future work.
Overall, while a coherent theoretical foundation of epsilon is missing, we consider the empirical results compelling.